Find the slope of the line passing through the points (5, 11) and (1, 3).
Answer by ilana(307)

Slope is (change in y)/(change in x), or (vertical change)/(horizontal change), or (y1-y2)/(x1-x2). Basically, subtract the y coordinates and subtract the x coordinates and divide. Just make sure your subtraction starts with the same (x,y) pair for x and y so you don't mess up with negatives.
I will start with (5,11) since those are both bigger than (1,3), so I will only have positives. 11-3=8, so the "change in y" is 8. 5-1=4, so the "change in x" is 4. Now divide: 8/4=2, so the slope is 2.